OpenStudy (anonymous):

The technical crew of the play needs shirts. Each shirt color matches the job the crew will be working on. Different color shirts have different designs and therefore cost different amounts. You will need:  6 red shirts, at $10.90 each for the lighting crew 4 yellow shirts at $ $10.50 each for the props crew  5 purple shirts at $12.50 each for the make-up crew You have $200.00 to spend on all the shirts. How many green shirts can you purchase for the ushers if those shirts cost $11.50 each? If you spend less than your budget, how much money do you have left over?

OpenStudy (texaschic101):

6(10.90) + 4(10.50) + 5(12.50) = 65.40 + 42 + 62.50 = 169.90 200 - 169.90 = 30.10 (this is what you have left after the purchase of the red,yellow, and purple shirts). 30.10 - 2(11.50) = 30.10 - 23 = 7.10 He can then buy 2 green shirts and have $7.10 change left.